# On-line conformalized neural networks ensembles for probabilistic forecasting of day-ahead electricity prices

### v1

@@ 2.4 Prediction quantiles estimation and DE combination
The output layers of the NNs and the loss functions are specialized accordingly, i.e., providing the predicted quantiles averaged by the pinball score in the former, parameterizing the target distribution then passed through a negative log-likelihood in the latter.
@@
As the state-of-art Distributional NNs for PEPF is based on multi-step feed-forward maps, we follow the same structural design in this work. This enables a proper comparison of the proposed PEPF technique under coherent settings (see Section 3). The investigation of further NN architectural forms is left to future studies. Such NN configuration is labelled DNN hereafter (i.e., Deep Neural Network), following the naming convention in [30] [4].
@@
The last design stage regards the definition of the approach to combine the neural network components within the DE. To this end, a wide set of methods have been studied in the time-series forecasting literature, from uniform marginal aggregations, to complex conditional weighting (see [46] for a recent detailed review). Still, the uniform aggregation (i.e., ensemble combination considering equal weights) have been shown to provide a robust alternative difficult to beat in practical settings. Similar observations have been reported under both point [4] and probabilistic [30] DNN-based settings in EPF studies. Moreover, it constitute a widely adopted approach in the broader Deep Learning literature devoted to DE (see e.g., [47]). Therefore, we exploit a uniform ensemble combination for the present study, leaving the investigation of alternative techniques to future extensions. For point predictors (e.g., in the case of QRA and base CP in the next section), this results in a simple average between the DNN components. For Distributional NNs, both probability aggregation and vincentization techniques can be considered. We resort to the latter since it typically leads to sharper bounds, beside marginalizing the different local minimizers reached by the training algorithm [46]. Formally, this is computed for each prediction stage and quantile as:
$q^{h}_{\gamma}(x_{i})=\sum_{j=1}^{n_{e}}\tilde{q}_{\gamma}^{h(j)}(x_{i}),\;\forall{\gamma}\in\Gamma,\;\forall{h}\in[1,...,H]$ (16)
@@ 3.1 Case studies
As case studies, we focused on both the German market (GE) from [30] and the different bidding zones constituting the Italian day-ahead markets (namely NOR, CNOR, CSOU, SOU, SARD, SICI) made available by [39], providing a compelling setup for comparative evaluations under heterogeneous conditions. The exogenous set employed for the GE market includes the day-ahead load and renewable generation forecast as well as the most recent gas closing prices. The Italian datasets comprise the hourly load and wind generation predictions.
@@
As input features, we included the subset selected with frequency 100% during the experiments performed in [30]. To summarize, we adopted: the price values over the previous 2 days $t$-1, $t$-2, i.e., 48 lags; the day-ahead load forecasts available at time $t$; the renewable energy sources predictions for the target date $t$ as well as the previous day $t$-1; the most recent closing gas price available, i.e. at $t$-2; the weekday encoding. The motivation behind such choice is twofold. On the one hand, we aim to explore a group similar to the one employed in the previous work. On the other hand, we target the assessment of the different PEPF models under consistent input variables. [^1] In [30], the authors reported sensible variations both in terms of the hyper-parameters chosen (including NNs input features selection via indicator variables) and the consequent test performances (see [30] for further details). Still, a deepen dataset analysis and feature selection may lead to better average performances. We leave such investigation to future extensions.
@@
$c_{t}=\left[\text{sin}\left(2\pi d_{t}/6\right),\ \ \text{cos}\left(2\pi d_{t}/6\right)\right]$ (17)
@@ 3.2 Benchmarks
Moreover, we implement CP techniques on both normal distributions and conventional absolute residuals (see [31]) to assess the benefit of the more flexible quantile-level corrections introduced. The former has been implemented following the same setup of the other distributional NN models to achieve a fair comparison. Hence, each NN in the ensemble parameterizes the mean and scale of the Gaussian distribution, followed by quantile vincentization. Although normalized conformity scores may be alternatively computed through the mean and variance of a mixture with uniform weights, probabilistic aggregation has been shown to perform worse than quantile ensembling (see [30], [48])  [^3] The exploration of conformal inference on further forecasting combination techniques represent an interesting direction of future research.. The CP on absolute residuals is obtained from the average predictions of NNs trained by minimizing the Mean Absolute Error.
@@ 3.3 Experimental setup
The experiments have been performed by leveraging Tensorflow, including the Tensorflow Probability library which provides several utilities to implement distributional NNs. [^4] To setup the backbone DE framework, the definition of the configurations follows the one used in the analysis reported in [30] for the baseline distributional and quantile regression based methods.
@@ 3.4 Evaluation metrics and tests
To evaluate the probabilistic forecasting performances achieved on the test sets, we follow the common practices in the PEPF field (see e.g. [1], [30] and references therein for further details). Calibration is first analyzed by means of the Kupiec test (with significance level 0.05) for unconditional coverage (see Section 3.4.1), on both extreme and central PIs, for each day-ahead hour. Then, the average Pinball loss across the deciles (see Section 2) and the Winkler’s score (see Section 3.4.2) for the related miscoverage levels are employed as proper scoring rules. It is worth noting that the average Pinball score, computed over the distribution percentiles, is commonly employed in PEPF as a cheap discrete approximation to the continuous ranked probability score (CRPS) (see e.g. [1]) Statistical significance is evaluated via the multivariate Diebold and Mariano (DM) test (see Section 3.4.3) on the differences in the loss norms between competing models. Point prediction errors are evaluated beyond the probabilistic forecasting performances by means of the Mean Absolute Error (MAE):
@@ 3.4.2 Winkler’s score
The Winkler’s score (or interval score) is a proper scoring rule to assess probabilistic forecasts formed as Prediction Intervals (PI) at discrete coverage levels $1-\alpha$ (see [50, Section 5.9]). Formally, it is defined as:
@@ 3.5 Results analysis
Further insights are provided by the hourly Prediction Interval Coverage Probability (PICP) values, reported in Figure 9 and Figure 10 for $\alpha$=0.2 and $\alpha$=0.6 respectively, and the average PICPs for each coupled deciles in Table 2. To support graphical readability, we have structured three subplots aggregating: the benchmark methods in the left column; the different conformalized quantiles setups (CQ*) in the middle column; the on-line conformal PI control integration (OCQ*) in the right column. The QRA benchmark coverage is included in all subplots to provide a coherent base within the common y axis scale. The red dashed lines represent the target coverage value $1-\alpha$777e.g., $0.81-\alpha=$ 0.8 for $\alpha=0.2$.
@@
Beside being more reliable, the conformalized models have preserved stable and in some cases improved scores, as shown in Table 3. The multivariate DM tests performed on the Pinball and Winkler’s score are displayed in Figure 13 and Figure 13 respectively, where the coloured cells highlight the p-value for the difference between model’s predictions (see Section 3.4.3). The test set Winkler’s scores are computed for each of the PI1-α obtained from the approximated distribution deciles.
@@
To provide further insights on the different price behaviours addressed by the models among the regional markets, we plot in Figure 14 extracts from GE, NOR and SICI. The deciles predicted by the OCQR model are reported besides the target series. Clearly, the price in each zone assume specific shapes and values, which depends on the generation mix (e.g., share of wind power generation in the southern part of Italy) and the different power demands. We leave to [39] and references therein for further details on the specific characteristics of the different Italian bidding area, and to [10] for a recent detailed analysis of the German power market.
@@ 4 Conclusions and next developments
The goal of PEPF, as probabilistic forecasting in general, is to maximize sharpness while respecting calibration requirements. In fact, efficient but reliabile uncertainty quantifications are decisive in supporting decision making under stochastic conditions, as it is the case in modern day-ahead EPF applications. In this work, we addressed the limited hourly calibration of the state of the art neural network (NN) ensemble based methods for PEPF by exploiting a flexible conformal prediction (CP) framework. To this end, the quantiles approximated by the NNs at each prediction step have been incrementally conformalized over the test set following a daily retraining (i.e., recalibration). Both quantile regression and distributional NNs approaches have been exploited for estimating the conditional quantiles to be calibrated. A uniform vincentization technique has been employed to aggregate the quantiles produced by the underlying ensemble components, providing a simple but robust forecasting combination in practice. Besides, a post-hoc sorting operator has been included to achieve conditional quantile non-crossing. The asymmetric Conformalized Quantile Regression formulation, adjusted in a multi-horizon forecasting setup, have been deployed to compute the scores and the conformalized quantiles. This supports a proper correction of both upper and lower bounds at each stage over the prediciton horizon, leading to a more efficient local overage in the features conditioned PIs estimation. Such capability is particularly crucial to tackle the complex distributional patterns typically observed in PEPF applications, such as sensible heteroskedasticity, skewness and fat tails. Besides, it compensates potential quantile overfitting of the conventional pinball loss based neural network training, enhancing the asymptotic validity with finite samples coverage guarantees. To address the lack of robustness to failures of the samples exchangeability assumptions, we have exploited an adversarial setup. Quantiles tracking and coverage error integration stages have been included, aimed to compensate potential coverage degradation and systematic errors occurring over the recalibration window, e.g., due to sensible short-term drifts. The whole procedure have been started on a samples subset close to the first test date to acquire the initial bag of calibration scores, then proceeding in a rolling window fashion. Indeed, updating the ensemble components by including the most recent observations support probabilistic performances beside point accuracy, as the underlying model quality impacts the local reliability of the prediction intervals despite the marginal coverage guarantees provided by the CP framework. The ensemble aggregation provides a further mechanisms for this purpose, by getting rid of poor local minimizers in the network parameter space reached by single trainings.

### v2

@@ 2.4 Prediction quantiles estimation and DE combination
The output layers of the NNs and the loss functions are specialized accordingly, i.e. providing the predicted quantiles averaged by the pinball score in the former, parameterizing the target distribution then passed through a negative log-likelihood in the latter.
@@
As the state-of-art Distributional NN for PEPF is based on multi-step feed-forward maps, we follow the same structural design in this work. This enables a proper comparison of the proposed PEPF technique under coherent settings (see section 3). The investigation of further NN architectural forms is left to future studies. Such NN configuration is labelled DNN hereafter (i.e., Deep Neural Network), following the naming convention in [33] [4])
@@
The last design stage regards the definition of the approach to combine the neural network components within the DE. To this end, a wide set of methods have been studied in the time-series forecasting literature, from uniform marginal aggregations, to complex conditional weighting (see [50] for a recent detailed review). Still, the uniform aggregation (i.e., ensemble combination considering equal weights) have been shown to provide a robust alternative difficult to beat in practical settings. Similar observations have been reported under both point [4] and probabilistic [33] DNN-based settings EPF studies. Moreover, it constitute a widely adopted approach in the broader Deep Learning literature devoted to DE (see e.g., [51]) Therefore, we exploit a uniform ensemble combination for the present study, leaving the investigation of alternative techniques to future extensions. For point predictors (e.g., in the case of QRA and base CP in the next section), this results in a simple average between the DNN components. For Distributional NNs, both probability aggregation and vincentization techniques can be considered. We resort to the latter since it typically leads to sharper bounds, beside marginalizing the different local minimizers reached by the training algorithm [50]. Formally, this is computed for each prediction stage and quantile as:
$q^{h}_{\gamma}(x_{i})=\sum_{j=1}^{n_{e}}\tilde{q}_{{\gamma}}^{h\mathbf{(j)}}(x_{i}),\;\forall{\gamma}\in\Gamma,\;\forall{h}\in[1,...,H]$ (16)
@@ 3.1 Case studies
As case studies, we focused on both the German market (GE) from [33] and the different bidding zones constituting the Italian day-ahead markets (namely NOR, CNOR, CSOU, SOU, SARD, SICI) made available by [43], providing a compelling setup for comparative evaluations under heterogeneous conditions. The exogenous set for the GE market includes the day-ahead load and renewable generation forecast as well as the most recent gas closing prices. The Italian datasets comprise the hourly load and wind generation predictions.
@@
As input features, we included the subset selected with frequency 100% during the experiments performed in [33] To summarize, we adopted: the price values over the previous 2 days $t$-1, $t$-2, i.e., 48 lags; the day-ahead load forecasts available at time $t$; the renewable energy sources predictions for the target date $t$ as well as the previous day $t$-1; the most recent closing gas price available, i.e. at $t$-2; the weekday encoding. The motivation behind such choice is twofold. On the one hand, we aim to explore a group similar to the one employed in the previous work. On the other hand, we target the assessment of the different PEPF models under consistent input variables. [^4] In [33], the authors reported sensible variations both in terms of the hyper-parameters chosen (including NNs input features selection via indicator variables) and the consequent test performances (see [33] for further details). Still, a deepen dataset analysis and feature selection may lead to better average performances. We leave such investigation to future extensions.
@@
$c_{t}=\left[\text{sin}\left(2\pi d_{t}/7\right),\ \ \text{cos}\left(2\pi d_{t}/7\right)\right]$ (17)
@@ 3.2 Benchmarks
Moreover, we implement CP techniques on both normal distributions and conventional absolute residuals (see [34]) to assess the benefit of the more flexible quantile-level corrections introduced. The former has been implemented following the same setup of the other distributional NN models to achieve a fair comparison. Hence, each NN in the ensemble parameterizes the mean and scale of the Gaussian distribution, followed by quantile vincentization. Although normalized conformity scores may be alternatively computed through the mean and variance of a mixture with uniform weights, probabilistic aggregation has been shown to perform worse than quantile ensembling (see [33], [52]).  [^6] The exploration of conformal inference on further forecasting combination techniques represent an interesting direction of future research. The CP on absolute residuals is obtained from the average predictions of NNs trained by minimizing the Mean Absolute Error.
@@ 3.3 Experimental setup
The experiments have been performed by leveraging Tensorflow, including the Tensorflow Probability library which provides several utilities to implement distributional NNs. [^7] To setup the backbone DE framework, the description of the configurations follows the one used in the analysis reported in [33] for the baseline distributional and quantile regression based methods.
@@ 3.4 Evaluation metrics and tests
To evaluate the probabilistic forecasting performances achieved on the test sets, we follow the common practices in the PEPF field (see e.g. [1] and references therein for further details). Calibration is first analyzed by means of the Kupiec test (with significance level 0.05) for unconditional coverage, on both extreme and central PIs, for each day-ahead hour. Then, the average Pinball loss across the deciles (see section 2) and the Winkler’s score for the related miscoverage levels are employed as proper scoring rules. It is worth noting that the average Pinball score, computed over the distribution percentiles, is commonly employed in PEPF as a cheap discrete approximation to the continuous ranked probability score (CRPS) (see e.g. [1]) Statistical significance is evaluated via the multivariate Diebold and Mariano (DM) test on the differences in the loss norms between competing models. Point prediction errors are evaluated beyond the probabilistic forecasting performances by means of the Mean Absolute Error (MAE):
@@ 3.4.2 Winkler’s score
The Winkler’s score (or interval score) is a proper scoring rule to assess probabilistic forecasts formed as Prediction Intervals (PI) at discrete coverage levels $1-\alpha$ (see [54, section 5.9]). Formally, it is defined as:
@@ 3.5 Results analysis
Further insights are provided by the hourly Prediction Interval Coverage Probability (PICP) values, reported in Figure 9 and Figure 10 for $\alpha$=0.2 and $\alpha$=0.6 respectively, and the average PICPs for each coupled deciles in Table 2. To support graphical readability, we have structured three subplots aggregating: the benchmark methods in the left column; the different conformalized quantile regression setups (CQ*) in the middle column; the on-line conformal PI control integration (OCQ*) in the right column. The QRA benchmark coverage is included in all subplots to provide a coherent base within the common y axis scale. The red dashed lines represent the target coverage value $1-\alpha$101010e.g., $0.81-\alpha=$ 0.8 for $\alpha=0.2$.
@@
Beside being more reliable, the conformalized models have preserved stable and in some cases improved scores, as shown in Table 3. Furthermore, the multivariate DM tests performed on the Pinball and Winkler’s score are displayed in Figure 13 and Figure 13 respectively, where the coloured cells highlight the p-value for the difference between model’s predictions (see Section 3.4.3). The test set Winkler’s scores are computed for each of the PI1-α obtained from the approximated distribution deciles.
@@
To provide further insights on the different price behaviours addressed by the models among the regional markets, we plot in Figure 14 extracts from GE, NOR and SICI. The deciles predicted by the CQR model are reported besides the target series. Clearly, the price in each zone assume specific shapes and values, which depends on the generation mix (e.g., share of wind power generation in the southern part of Italy) and the different power demands. We leave to [43] and references therein for further details on the specific characteristics of the different Italian bidding area, and to [10] for a recent detailed analysis of the German power market.
@@ 4 Conclusions and next developments
The goal of PEPF, as probabilistic forecasting in general, is to maximize sharpness while respecting calibration requirements. In fact, efficient but reliabile uncertainty quantifications are decisive in supporting decision making under stochastic conditions, as it is the case in modern day-ahead EPF applications. In this work, we addressed the limited hourly calibration of the state of the art neural network (NN) ensemble based methods for PEPF by exploiting a flexible Conformal Prediction (CP) framework. To this end, the quantiles approximated by the NNs at each prediction step have been incrementally conformalized over the test set following a daily retraining (i.e., recalibration). Both quantile regression and distributional NNs approaches have been exploited for estimating the conditional quantiles to be calibrated. A uniform vincentization technique has been employed to aggregate the quantiles produced by the underlying ensemble components, providing a simple but robust forecasting combination in practice. Besides, a post-hoc sorting operator has been included to achieve conditional quantile non-crossing. The asymmetric Conformalized Quantile Regression formulation, adjusted in a multi-horizon forecasting setup, have been deployed to compute the scores and the conformalized quantiles. This supports a proper correction of both upper and lower bounds at each stage over the prediciton horizon, leading to a more efficient local overage in the features conditioned PIs estimation. Such capability is particularly crucial to tackle the complex distributional patterns typically observed in PEPF applications, such as sensible heteroskedasticity, skewness and fat tails. Besides, it compensates potential quantile overfitting of the conventional pinball loss based neural network training, enhancing the asymptotic validity with finite samples coverage guarantees. To address the lack of robustness to failures of the samples exchangeability assumptions, we have exploited an adversarial setup. Quantiles tracking and coverage error integration stages have been included, aimed to compensate potential coverage degradation and systematic errors occurring over the recalibration window, e.g., due to sensible short-term drifts. The whole procedure have been started on a samples subset close to the first test date to acquire the initial bag of calibration scores, then proceeding in a rolling window fashion. Indeed, updating the ensemble components by including the most recent observations support probabilistic performances beside point accuracy, as the underlying model quality impacts the local reliability of the prediction intervals despite the marginal coverage guarantees provided by the CP framework. Besides, the ensemble aggregation provides a further mechanisms for this purpose, by getting rid of poor local minimizers in the network parameter space reached by single trainings.